# SVDQuant: Absorbing Outliers by Low-Rank Components for 4-Bit Diffusion Models

### v2

@@ 6 Conclusion & Discussion
In this work, we introduce a novel 4-bit post-training quantization paradigm, SVDQuant, for diffusion models. It adopts a low-rank branch to absorb the outliers in both the weights and activations, easing the process of quantization. Our inference engine Nunchaku further fuses the low-rank and low-bit branch kernels, reducing memory usage and cutting off redundant data movement overhead. Extensive experiments demonstrate that SVDQuant preserves image quality. Nunchaku further achieves a 3.5× reduction in memory usage over the original 16-bit model and 3.0× speedup over the weight-only quantization on an NVIDIA RTX-4090 laptop. This advancement enables the efficient deployment of large-scale diffusion models on edge devices, unlocking broader potential for interactive AI applications.
Limitations. In this work, we do not report the speedups for our FP4 models. This is because we have no access to Blackwell GPUs, which natively support the precision and microscaling for group quantization. On Blackwell hardware, we anticipate greater speedups compared to our INT4 results on 4090 GPUs.
@@ Acknowledgments
Appendix A Missing Proofs
@@
$\leq$ $\frac{\sqrt{\text{size}({\bm{R}})}}{q_{\max}}\cdot\sigma\sqrt{2\log\left(\text{size}({\bm{R}})\right)},$ (8)
@@
Together, we have
$\mathbb{E}\left[\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}\right]$
$\leq$ $\frac{\sqrt{\text{size}({\bm{R}})}}{q_{\max}}\cdot\sigma\sqrt{2\log\left(\text{size}({\bm{R}})\right)}$
$\leq$ $\frac{\sqrt{\log\left(\text{size}({\bm{R}})\right)\pi}}{q_{\max}}\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right].$
Appendix B Additional Results
B.1 Quality Results
We report extra quantitative quality results with additional metrics in Table 2. Specifically, CLIP IQA (Wang et al., 2023) and CLIP Score (Hessel et al., 2021) assesses the image quality and text-image alignment with CLIP (Radford et al., 2021), respectively. Structural Similarity Index Measure (SSIM) is used to measure the luminance, contrast, and structure similarity of images produced by our 4-bit model against the original 16-bit model. We also visualize more qualitative comparsions in Figures 12, 13, 14, 15 and 16.
Table 2: Additional quantitative quality comparisons across different models. C.IQA means CLIP IQA, and C.SCR means CLIP Score.
@@ B.1 Quality Results
INT W4A4 Ours 0.927 26.6 0.602 0.952 26.1 0.519
FP W4A4 Ours 0.935 26.7 0.652 0.957 26.1 0.574
@@ B.2 LoRA Results
Figure 17: Additional LoRA results on FLUX.1-dev. When applying LoRAs, our INT4 model matches the image quality of the original BF16 model. See Appendix C for the detailed used text prompts.
Appendix C Text Prompts

### v3

@@ 6 Conclusion
In this work, we introduce a novel 4-bit post-training quantization paradigm, SVDQuant, for diffusion models. It adopts a low-rank branch to absorb the outliers in both the weights and activations, easing the process of quantization. Our inference engine Nunchaku further fuses the low-rank and low-bit branch kernels, reducing memory usage and cutting off redundant data movement overhead. Extensive experiments demonstrate that SVDQuant preserves image quality. Nunchaku further achieves a 3.5× reduction in memory usage over the original 16-bit model and 3.0× speedup over the W4A16 on an NVIDIA RTX 4090 and 5090 GPUs. This advancement enables the efficient deployment of large-scale diffusion models on edge devices, unlocking broader potential for interactive AI applications.
@@ Acknowledgments
Changelog
V1 Initial preprint release.
V2 Fix a typo.
V3 ICLR 2025 camera-ready version. Upgrade the models by combining SVDQuant and GPTQ. Update NVFP4 and SANA results.
Appendix A Proofs
@@
Proposition 4.1.
The quantization error $E({\bm{X}},{\bm{W}})=\left\lVert{\bm{X}}{\bm{W}}-Q({\bm{X}})Q({\bm{W}})\right\rVert_{F}$ in Equation 3 can be decomposed as follows:
@@
Proposition 4.2.
For any tensor ${\bm{R}}$ and quantization method described in Equation 1 as $Q({\bm{R}})=s_{\bm{R}}\cdot{\bm{Q}}_{\bm{R}}$. Assuming the elements of ${\bm{R}}$ follow a distribution that satisfies the following regularity condition: There exists a constant $c$ such that
$\mathbb{E}\left[\max(|{\bm{R}}|)\right]\leq c\cdot\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right].$ (10)
Then, we have
$\mathbb{E}\left[\left\lVert{\bm{R}}-Q({\bm{R}})\right\rVert_{F}\right]\leq\frac{c\sqrt{\text{size}({\bm{R}})}}{q_{\max}}\cdot\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right]$ (11)
where $\text{size}({\bm{R}})$ denotes the number of elements in ${\bm{R}}$. Especially if the elements of ${\bm{R}}$ follow a normal distribution, Equation 10 holds for $c=\sqrt{\frac{\log\left(\text{size}({\bm{R}})\right)\pi}{\text{size}({\bm{R}})}}$.
@@
$\leq$ $\frac{c\sqrt{\text{size}({\bm{R}})}}{q_{\max}}\cdot\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right]$
Especially, if the elements of ${\bm{R}}$ follows a normal distribution, we have
$\mathbb{E}\left[\max(|{\bm{R}}|)\right]\leq\sigma\sqrt{2\log\left(\text{size}({\bm{R}})\right)}$ (12)
@@
Together, we have that for a normal distribution,
$\mathbb{E}\left[\max(|{\bm{R}}|)\right]$
$\leq$ $\sigma\sqrt{2\log\left(\text{size}({\bm{R}})\right)}$
$\leq$ $\sqrt{\frac{\log\left(\text{size}({\bm{R}})\right)\pi}{\text{size}({\bm{R}})}}\mathbb{E}\left[\left\lVert{\bm{R}}\right\rVert_{F}\right].$
In other words, Equation 10 holds for $c=\sqrt{\frac{\log\left(\text{size}({\bm{R}})\right)\pi}{\text{size}({\bm{R}})}}$. ∎
Appendix B Benchmark Models
We benchmark our methods using the following six text-to-image models:
  • FLUX.1 (Black-Forest-Labs, 2024) is the SoTA open-sourced DiT-based diffusion model. It consists of 19 joint attention blocks (Esser et al., 2024) and 38 parallel attention blocks (Dehghani et al., 2023), totaling 12B parameters. We evaluate both the 50-step guidance-distilled (FLUX.1-dev) and 4-step timestep-distilled (FLUX.1-schnell) variants.
  • SANA (Xie et al., 2025) is a 1.6B DiT model. It utilizes a 32× compression autoencoder (Chen et al., 2025) and replaces Softmax attention with linear attention to accelerate image generation.
  • Stable Diffusion XL (SDXL) is a widely-used UNet-based model with 2.6B parameters (Podell et al., 2024). It predicts noise with three resolution scales. The highest-resolution stage is processed entirely by ResBlocks (He et al., 2016), while the other two stages jointly use ResBlocks and attention layers. Like PixArt-$\Sigma$, SDXL uses cross-attention layers for text conditioning. We evaluate it in the 30-step setting, along with its 4-step distilled variant, SDXL-Turbo (Sauer et al., 2024).
Appendix C Benchmark Datasets
To assess the generalization capability of our method, we adopt two distinct prompt sets with varying styles for benchmarking:
  • MJHQ-30K (Li et al., 2024a) consists of 30K samples from Midjourney with 10 common categories, 3K samples each. We randomly select 5K prompts from this dataset to evaluate model performance on artistic image generation.
  • Densely Captioned Images (DCI) (Urbanek et al., 2024) is a dataset containing $\sim$8K images with detailed human-annotated captions, averaging over 1,000 words. For our experiments, we use the summarized version (sDCI), where captions are condensed to 77 tokens using large language models (LLMs) to accommodate diffusion models. Similarly, we randomly sample 5K prompts for realistic image generation.
Appendix D Implementation Details
For the 8-bit setting, we use per-token dynamic activation quantization and per-channel weight quantization with a low-rank branch of rank 16. For the 4-bit setting, we adopt per-group symmetric quantization for both activations and weights, along with a low-rank branch of rank 32. INT4 quantization uses a group size of 64 with 16-bit scales. We use NVFP4 for FP4 quantization, which has native hardware support of group size of 16 with FP8 scales on Blackwell GPUs (NVIDIA Corporation, 2025). We use GPTQ (Frantar et al., 2023) to quantize the residual weights. For FLUX.1 models, the inputs of linear layers in adaptive normalization are kept in 16 bits (i.e., W4A16). For other models, key and value projections in the cross-attention are retained at 16 bits since their latency only covers less than 5% of total runtime.
The smoothing factor $\lambda\in\mathbb{R}^{m}$ is a per-channel vector whose $i$-th element is computed as $\lambda_{i}=\max(|{\bm{X}}_{:,i}|)^{\alpha}/\max(|{\bm{W}}_{i,:}|)^{1-\alpha}$ following SmoothQuant (Xiao et al., 2023) Here, ${\bm{X}}\in\mathbb{R}^{b\times m}$ and ${\bm{W}}\in\mathbb{R}^{m\times n}$. It is decided offline by searching for the best migration strength $\alpha$ for each layer to minimize the layer output mean squared error (MSE) after SVD on the calibration dataset.
Appendix E Additional Results
E.1 Visual Quality Results
We report extra quantitative quality results with additional metrics in Table 2. Specifically, CLIP IQA (Wang et al., 2023b) and CLIP Score (Hessel et al., 2021) assesses the image quality and text-image alignment with CLIP (Radford et al., 2021), respectively. Structural Similarity Index Measure (SSIM) is used to measure the luminance, contrast, and structure similarity of images produced by our 4-bit model against the original 16-bit model. We also visualize more qualitative comparsions in Figures 11, 12, 13, 14 and 15.
Table 2: Additional quantitative quality comparisons across different models. RTN stands for round-to-nearest. C.IQA means CLIP IQA, and C.SCR means CLIP Score.
@@ E.1 Visual Quality Results
INT W4A4 Ours 0.926 26.6 0.655 0.948 26.1 0.577
NVFP W4A4 Ours 0.938 26.7 0.692 0.956 26.1 0.618
BF16 0.934 26.8 0.958 26.4
SANA INT W4A4 RTN 0.915 26.9 0.604 0.943 26.4 0.538
-1.6B INT W4A4 Ours 0.926 26.9 0.710 0.951 26.4 0.649
(20 Steps) NVFP W4A4 RTN 0.929 26.8 0.694 0.953 26.4 0.626
NVFP W4A4 Ours 0.932 26.9 0.755 0.955 26.4 0.701
@@ E.2 LoRA Results
Figure 16: Additional LoRA results on FLUX.1-dev. When applying LoRAs, our INT4 model matches the image quality of the original BF16 model. See Appendix F for the detailed used text prompts.
E.3 Additional Ablation of SVDQuant
Table 3: Quantitative comparisons of different SVDQuant settings on MJHQ. NVFP4 outperforms INT4. SVDQuant leverages a low-rank branch to ease quantization, significantly enhancing image quality. It can further apply GPTQ to quantize the weight residual, further improving quality.
Model Precision Low-rank Branch GPTQ Image Reward ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$)
BF16 0.953
INT4 0.908 0.322 18.5
0.933 0.297 19.1
0.926 0.256 20.1
FLUX.1-dev 0.935 0.223 21.0
NVFP4 0.928 0.244 20.3
0.936 0.204 21.5
0.935 0.223 20.9
0.937 0.208 21.4
BF16 0.968
INT4 0.962 0.345 16.3
0.962 0.317 16.8
0.957 0.289 17.6
FLUX.1-schnell 0.951 0.258 18.3
NVFP4 0.957 0.280 17.5
0.956 0.247 18.5
0.968 0.247 18.4
0.968 0.227 19.0
BF16 0.944
INT4 -1.226 0.762 9.1
-0.902 0.763 9.9
0.858 0.356 17.0
PixArt-$\Sigma$ 0.878 0.323 17.6
NVFP4 0.660 0.517 14.8
0.696 0.480 15.6
0.945 0.290 18.0
0.940 0.271 18.5
BF16 0.952
INT4 0.894 0.339 15.3
0.881 0.288 16.4
0.922 0.234 17.4
SANA-1.6B 0.935 0.220 17.8
NVFP4 0.932 0.237 17.3
0.927 0.202 18.3
0.957 0.188 18.7
0.955 0.177 19.0
In Table 3, we provide additional quantitative ablation results of SVDQuant on the MJHQ prompt set (Li et al., 2024a). Across all models, NVFP4 outperforms INT4 due to its native support for smaller microscaling group sizes on Blackwell. SVDQuant leverages a low-rank branch to absorb outliers, significantly enhancing image quality in all settings. Additionally, it can incorporate GPTQ (Frantar et al., 2023) instead of round-to-nearest for weight quantization, further improving quality in most cases. Notably, combining SVDQuant with NVFP4 precision achieves the best results, reaching a PSNR of 21.5 on FLUX.1-dev, closely matching the image quality of the original 16-bit model. In Figure 17, we provide qualitative comparisons across different precision settings.
Figure 17: Qualitative comparisons of different precisions on MJHQ. NVFP4+SVDQuant yields the highest image fidelity.
E.4 Latency Results
In Table 4, we compare FLUX latency on a laptop-level 4090 GPU across different precisions. Compared to INT8, 4-bit quantization delivers a 1.3×1.3$\times$1.3 × speedup. However, without optimization, SVDQuant incurs an 18% overhead due to the low-rank branch. By eliminating redundant memory access, Nunchaku achieves latency comparable to naive INT4.
Table 4: Single-step latency comparisons of FLUX on a desktop-level 4090 GPU.
Method BF16 INT8 Naïve INT4 SVDQuant SVDQuant +Nunchaku
Latency (ms) 657 282 212 250 218
E.5 Trade-off of Increasing Rank
Figure 18 presents the results of different rank $r$ in SVDQuant on PixArt-$\Sigma$. Increasing the rank from 16 to 64 significantly enhances image quality but increases parameter and latency overhead. In our experiments, we select a rank of 32, which offers a decent quality with minor overhead.
E.6 Trade-off between Quality and Bitwidth
We evaluate LPIPS across different bitwidths for various quantization methods on PixArt-$\Sigma$ and FLUX.1-schnell using the MJHQ dataset in Figure 19, with weights and activations sharing the same bitwidth. Following the convention (Xiao et al., 2023; Lin et al., 2024; 2025; Li et al., 2023a; Zhao et al., 2024d; Dettmers et al., 2022), for bitwidths above 4, we apply per-channel quantization; for 4 or below, we use per-group quantization (group size 64). SVDQuant consistently outperforms naive quantization and SmoothQuant. Notably, on PixArt–$\Sigma$ and FLUX.1-schnell, our 4-bit results match 7-bit and 6-bit naive quantization, respectively.
Our SVDQuant can still generate images in the 3-bit settings on both PixArt-$\Sigma$ and FLUX.1-schnell, performing much better than SmoothQuant. Below this precision (e.g., W2A4 or W4A2), SVDQuant cannot produce images either since 2-bit symmetric quantization is essentially a ternary quantization. Prior work (Ma et al., 2024a; Wang et al., 2023a) has shown that ternary neural networks require quantization-aware training even for weight-only quantization to adapt the weights and activations to the low-bit distribution.
Figure 19: LPIPS of different quantization methods on PixArt-$\Sigma$ and FLUX.1-schnell across different bitwidths.
Appendix F Text Prompts